OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2

OpenStudy (amistre64):

given a point (Px,Py) and a slope m we can write: y-Py = m(x-Px) y = mx -mPx +Py ; where (-mPx +Py) forms a constant known as the y intercepts

OpenStudy (amistre64):

y = 2x -2(3) -2 ; and simplify